OpenStudy (anonymous):

HELP NOW PLEASE!!! Evaluate the expression Quantity 3a - 5 + 3. Match the variable pair with the resulting value of the expression. a=3 b=7

OpenStudy (cggurumanjunath):

3(3)-5+3=9-5+3=4+3=7

OpenStudy (anonymous):

and if a were -3 and b were -7

OpenStudy (cggurumanjunath):

3(-3)-5+3=-9-2=-11

OpenStudy (anonymous):

it can only be -17, 25 or -35